Silven Industries, which manufactures and sells a highly successful line of summer lotions and insect repellents, has decided to diversify in order to stabilize sales throughout the year. A natural area for the company to consider is the production of winter lotions and creams to prevent dry and chapped skin.

After considerable research, a winter products line has been developed. However, Silven’s president has decided to introduce only one of the new products for this coming winter. If the product is a success, further expansion in future years will be initiated.

The product selected (called Chap-Off) is a lip balm that will be sold in a lipstick-type tube. The product will be sold to wholesalers in boxes of 8 tubes for $6.00 per box. Because of excess capacity, no additional fixed manufacturing overhead costs will be incurred to produce the product. However, a $90,000 charge for fixed manufacturing overhead will be absorbed by the product under the company’s absorption costing system.

Using the estimated sales and production of 100,000 boxes of Chap-Off, the Accounting Department has developed the following cost per box:

Direct materials

$

2.20

Direct labor

1.00

Manufacturing overhead

1.50

Total cost

$

4.70

The costs above include costs for producing both the lip balm and the tube that contains it. As an alternative to making the tubes, Silven has approached a supplier to discuss the possibility of purchasing the tubes for Chap-Off. The purchase price of the empty tubes from the supplier would be $0.70 per box of 8 tubes. If Silven Industries accepts the purchase proposal, direct labor and variable manufacturing overhead costs per box of Chap-Off would be reduced by 10% and direct materials costs would be reduced by 20%.

Required:

1a. Calculate the total variable cost of producing one box of Chap-Off? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

1b. Assume that the tubes for the Chap-Off are purchased from the outside supplier, calculate the total variable cost of producing one box of Chap-Off? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

1c. Should Silven Industries make or buy the tubes?

Make

Buy

   

2. What would be the maximum purchase price acceptable to Silven Industries? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

     

3. Instead of sales of 100,000 boxes, revised estimates show a sales volume of 110,000 boxes. At this new volume, additional equipment must be acquired to manufacture the tubes at an annual rental of $26,000. Assume that the outside supplier will not accept an order for less than 124,000 boxes.

a. Calculate the total relevant cost of making 124,000 boxes and total relevant cost of buying 124,000 boxes. (Do not round intermediate calculations.)

b. Based on the above calculations, should Silven Industries make or buy the boxes?

Make

Buy

  

4. Refer to the data in (3) above. Assume that the outside supplier will accept an order of any size for the tubes at $0.70 per box. Which of these is the best alternative?

Make all 124,000 boxes

Buy all 124,000 boxes

Make 100,000 boxes and buy 24,000 boxes

Make 26,000 boxes and buy 26,000 boxes
Answer #1
1-a. Total variable cost per box:
$
Direct materials 2.2
Direct labor 1
Variable Manufacturing overhead 0.6
(Note:1)
Total 3.8
Note:1
$
Total manufacturing overhead 1.5
Less: Fixed manufacturing overhead
(Total fixed manufacturing overhead/Estimated production)
(90000/100000) 0.9
Variable Manufacturing overhead 0.6
1-b. If purchased from outside supplier,
Total variable cost per box:
$
Direct materials 2.2*(100%-20%) 1.76
Direct labor 1*(100%-10%) 0.9
Variable Manufacturing overhead 0.6*(100%-10%) 0.54
Cost of tube purchased 0.7
Total 3.9
1-c Make
Since variable cost per box is less in this case
2 Maximum purchase price=Variable cost of producing a tube+Change in variable cost due to buying from outside
Change in variable cost due to buying from outside:
$
Direct materials 2.2*20% 0.44
Direct labor 1*10% 0.1
Variable Manufacturing overhead 0.6*10% 0.06
Total 0.6
Maximum purchase price=3.8+0.6=$ 4.4 per box
3-a. Total relevant cost of making:
$
Total variable manufacturing cost (Refer 1-a) (124000*3.8) 471200
Annual rental 26000
Total relevant cost of making 497200
Total relevant cost of buying=124000*variable cost per box if purchased from outside=124000*3.9=$ 483600
3-b. Buy
Since total relevant cost is less in this case
4 Total cost of making 124000 boxes=$ 497200
Total cost of buying 124000 boxes=$ 483600
Total cost of making 100000 boxes and buy 24000 boxes=(100000*3.8)+(24000*3.9)=$ 473600
Make 26000 boxes and buy 26000 boxes will not meet 124000 boxes.Hence,avoid this option
Hence,best option is Make 100000 boxes and buy 24000 boxes since it's total cost is less compared to other options